Equipment acquired at a cost of $75,000 has an estimated residual value of $5,000 and an estimated useful life of 10 years. It was placed in service on October 1 of the current fiscal year, which ends on December 31.

a. Determine the depreciation for the current fiscal year and for the following fiscal year by the straight-line method.

b. Determine the depreciation for the current fiscal year and the following fiscal year by the double-declining-balance method.
Answer #1

Answer a.

Depreciation under straight line method
1st year $    1,750
2nd year $    7,000
Answer b.
Depreciation under double declining method
1st year $    3,750
2nd year $ 14,250

working for the above answer

Straight line method
Capitalized Cost $                75,000
Life of Assets 10 Years
Residual value $                  5,000
Depreciation = (Cost-Residual Value)/Life years =(75000-5000)/10
=70000/10= $                  7,000
Depreciation for 1st year from 1st Oct to 31st December
=7000*3/12 = $                  1,750
Depreciation for 2nd year $                  7,000
Double Declining Balance Method
Capitalized Cost $                75,000
Life of Truck 10 Years
Residual value $                  5,000
Rate of Depreciation under DDB = 2*100/10 20%
First yea Depreciation = 20% of $75000 for 3 months =
=75000 x 20% x 3/12 $                  3,750
Book value in the beginning of 2nd year $                71,250
Depreciation for 2nd year $                14,250
=20% of ($75000-$3750) =

Hit Thumbs up if satisfied

Have any query mention in comment section please

Thank you